If the inventor charges $4 per unit, then her profit for producing and selling x units is given by the function
P(x) = 2.25x -7000
(a) What is her profit if she sells 2000 units?
(b) What is her profit if she sells 5000 units?
(c) What is the break-even point for sales?

Answer by checkley71(8403)

P(X)=2.25X-7000
P(2000)=2.25*2000-7000
P(2000)=4500-7000
P(2000)=-2500 LOSS FOR 200 UNITS.
P(5000)=2.25*5000-7000
P(5000)=11250-7000
P(5000)=4250 PROFIT.
0=2.25X-7000
2.25X=7000
X=7000/2.25
X=3111 UNITS ARE REQUIRED TO BREAK EVEN.